The ground state wave function of a particle in a 3D potential is given by ¥=Ae^{ar^3}, where A is the normalisation constant. The value of A is given by
1
Expert's answer
2019-05-13T10:42:58-0400

Let a<0. Normalization condition:


"\int \psi^2 r^2 dr=1"

(integral limits: from 0 to infinity ).

"\int A^2 e^{2ar^3} r^2 dr=(A^2/6a)\int e^{2ar^3} d(2ar^3)=(A^2/6a)e^{2ar^3}"

Substituting the limits of the integral and equating the expression to 1:


"-A^2\/(6a)=1"

"A^2=-6a"

"A1=\sqrt{-6a}"